# Weyl law for 1-cycles

### v1

@@ 3.2. Main Theorems
In this subsection, we state the approximation results by $\delta$-localized families in the discrete and continuous setting. These are crucial in the proof of the Parametric Isoperimetric Inequality discussed later in this article and also to obtain the Parametric Coarea Inequality in [24].
@@
Given $10\leq k\leq n-$ 1 and $p,N\in\mathbb{N}$ there exist natural numbers $K(N,k,p)$, $n(k,p)$ and positive constants $c(k,p,M)$, $\varepsilon_{0}(k,p,M)$ and $\delta_{0}(k,p,M)$ so that the following is true. Let $\varepsilon\leq\varepsilon_{0}$, $\delta\leq\delta_{0}$ and let $F:X_{0}\to\mathcal{Z}_{k}(M)$ be a discrete $(N,\delta)$-localized family of cycles with $\mathcal{F}(F(x))\leq\varepsilon$ for every $x\in X_{0}$ and $\dim(X)=p$. Then there exists a $(K(N,k,p),K(N,k,p)\delta)$-localized family $\tau:X(q)_{0}\to\mathcal{I}_{k+1}(M)$ such that $\partial\tau(x)=R^{q}F(x)$ for $q=q(k,p,M,\delta)$ and $\operatorname{\mathbf{M}}(\tau(x))\leq c(k,p,M)\frac{\varepsilon}{\delta^{n(k,p)}}$.
Roughly speaking, the previous theorem tells us that if we have a discrete family of $k$-cycles $F:X_{0}\to\mathcal{Z}_{k}(M)$ which has flat norm at most $\varepsilon$ at every point and is $\delta$-localized for $\varepsilon$ and $\delta$ small enough, then we can construct a discrete filling with similar properties. To be precise, the filling is a $\delta^{\prime}$-localized family of $(k+1)$-chains $\tau$ such that up to refinement $\partial\tau(x)=F(x)$, $\delta^{\prime}$ is controlled in terms of $\delta$ and $\operatorname{\mathbf{M}}(\tau(x))$ is controlled in terms of $\varepsilon$ and $\delta$. Theorem 3.12 is proved in subsection 3.3 and is used to prove the following result.
@@
There exist constants $\varepsilon_{0}(k,p,M)>0$, $\delta_{0}(k,p,M)>0$, $C(k,p,M)>0$, $n(k,p)\in\mathbb{N}$ and $N(k,p)\in\mathbb{N}$ such that the following is true. Let $F:X_{0}\to\mathcal{Z}_{k}(M)$ be an $\varepsilon$-fine family, $0<\varepsilon\leq\varepsilon_{0}$, $10\leq k\leq n-$ 1, $\dim(X)=p$. Let $0<\delta<\delta_{0}$. Then there exists a discrete family $F^{\prime}:X(q)_{0}\to\mathcal{Z}_{k}(M)$ for $q=q(k,p,M,\delta)$ such that
@@
There exist constants $\varepsilon_{0}(k,p,M)>0$, $\delta_{0}(k,p,M)>0$, $C(k,p,M)$, $n(k,p)\in\mathbb{N}$ and $N(k,p)\in\mathbb{N}$ such that the following is true. Let $G:X_{0}\to\mathcal{I}_{k+1}(M)$ be an $\varepsilon$-fine family, $0<\varepsilon\leq\varepsilon_{0}$, $10\leq k\leq n-$ 1, $\dim(X)=p$ and denote $F=\partial G$. Let $0<\delta<\delta_{0}$. Then there exists a discrete family $G^{\prime}:X(q)_{0}\to\mathcal{I}_{k+1}(M)$ for $q=q(k,p,M,\delta)$ such that if we denote $F^{\prime}=\partial G^{\prime}$ then
@@
Let $G:X\to\mathcal{I}_{k+1}(M)$ be a continuous map without concentration of mass. Denote $F=\partial G$ and let $\varepsilon,\delta>0$. Then there exists $q,q^{\prime}\in\mathbb{N}$ and a continuous map $G^{\prime}:X\to\mathcal{I}_{k+1}(M)$ without concentration of mass such that
@@
3.3. Filling small families of cycles
@@
Given $10\leq k\leq n-$ 1 and $p,N\in\mathbb{N}$ there exist natural numbers $K(N,k,p)$, $n(k,p)$ and positive constants $c(k,p)$, $\varepsilon_{0}(k,p,M)$ and $\delta_{0}(k,p,M)$ so that the following is true. Let $\varepsilon\leq\varepsilon_{0}$, $\delta\leq\delta_{0}$ and let $F:X_{0}\to\mathcal{Z}_{k}(M)$ be a discrete $(N,\delta)$-localized family of cycles with $\mathcal{F}(F(x))\leq\varepsilon$ for every $x\in X_{0}$ and $\dim(X)=p$. Then there exists a $(K(N,k,p),K(N,k,p)\delta)$-localized family $\tau:X(q)_{0}\to\mathcal{I}_{k+1}(M)$ such that $\partial\tau(x)=R^{q}F(x)$ for $q=q(k,p,M,\delta)$ and $\operatorname{\mathbf{M}}(\tau(x))\leq c(k,p)(\frac{L}{\delta})^{n(k,p)}\varepsilon$.
In order to prove the proposition, we need to introduce some definitions and do some constructions first. Let $r<\operatorname{inj}(M,g)$ and consider a collection of $L$ points $x_{1},...,x_{L}\in M$ such that $\{B(x_{l},r):1\leq l\leq L\}$ covers $M$ and $L\leq c(n)\frac{\operatorname{Vol}(M,g)}{r^{n}}$ for some universal constant $c(n)$ depending only on $n$. For this section, we will consider be $r=\delta$ and the corresponding $L$ is the one that appears in the statement of Proposition 3.17. In the next sections, we will choose other values of $r$, being that the reason why we do the following constructions for arbitrary $r<\operatorname{inj}(M,g)$.
@@
$\operatorname{\mathbf{M}}(\tau\llcorner\partial B(x_{l},r_{l}(\tau))\leq\frac{\operatorname{\mathbf{M}}(\tau)}{r}.$
@@
We introduce the following definitions which will be used in the rest of the section.
@@
$d^{q}(x,y)$ equals the minimal number of vertices in $X(q)_{0}$ that we must visit if we want to travel from $x$ to $y$ by jumps between vertices in $X(q)_{0}$ differing in only one coordinate. When $q$ is clear (because we are considering $x,y\in X(q)_{0}$ for some specific value of $q$) we may just denote $d^{q}$ by $d$.
@@
$d_{l}(\tau)-d_{l}(\tau^{\prime})$ $=\tau-\tau^{\prime}-[\tau\llcorner\bigcup_{m=1}^{l}B(x_{m},r_{m})-\tau^{\prime}\llcorner\bigcup_{m=1}^{l}B(x_{m},r^{\prime}_{m})]$
$=\tau-\tau^{\prime}-[\tau\llcorner\bigcup_{m=1}^{l}B(x_{m},r_{m})\Delta B(x_{m},r^{\prime}_{m})+(\tau-\tau^{\prime})\llcorner\bigcup_{m=1}^{l}B(x_{m},r_{m}^{\prime})]$
which implies that
@@
First we prove $(N,\delta)$-localization. Let $y_{1},y_{2}$ be vertices of the same cell $E$ in $C(3q)$. If $y_{2},y_{3}\in\operatorname{Center}(C)$ the result follows so we can assume $l(y_{1}),l(y_{2})\leq q$. Let $E^{\prime}$ be a $(j-1)$-face of $\partial C(q)$ such that $\Xi(E)\subseteq E^{\prime}$. Then without loss of generality, there exists another vertex $y_{3}$ of $E$ such that $\Xi(y_{2})=\Xi(y_{3})$, $l(y_{2})=l(y_{3})+1$, $l(y_{1})=l(y_{3})$ and $\Xi(y_{1}),\Xi(y_{3})$ belong to $E^{\prime}$. By Lemma 3.22 applied to the chains $\tau=\tau(\Xi(y_{1}))$, $\tau^{\prime}=\tau(\Xi(y_{3}))$, the number $l=l(y_{1})=l(y_{3})$ and the $(N,\delta)$-admissible family $\{U_{i}\}_{i\in I_{E^{\prime}}}$ with $U_{i}=B(y_{i},s_{i})$ it holds
@@
For each cell $C$ of $X$, let $\{U_{i}^{C}\}_{i\in I_{C}}$ be the $\delta$-admissible family associated to $C$ and let $U_{C}=\bigcup_{i\in I_{C}}U_{i}^{C}$. We proceed by induction on $n-k$, being $k=n-1$ the base case.
Base case 1. Let $k=n-1$. For each $x\in X_{0}$, as $\mathcal{F}(F(x))\leq\varepsilon$ there exists an $n$-chain $\tau(x)$ such that $\partial\tau(x)=F(x)$ and $\operatorname{\mathbf{M}}(\tau(x))\leq\varepsilon$. Consider the corresponding family $\tau:X_{0}\to\mathcal{I}_{n}(M)$. We claim that if $C$ is a cell of $X$ and $x,y\in X_{0}\cap C$ then $\tau(x)-\tau(y)$ is supported in $U_{C}$. Indeed, as $\operatorname{spt}(F(x)-F(y))\subseteq U_{C}$, there exists $\tau_{xy}\in\mathcal{I}_{n}(U_{C})$ such that $\partial\tau_{xy}=F(x)-F(y)$ and $\operatorname{\mathbf{M}}(\tau_{xy})\leq 2\varepsilon$. Therefore $\eta_{xy}=\tau(x)-\tau(y)-\tau_{xy}$ is an $n$-cycle on $M$ of mass at most $4\varepsilon$. Thus if we set $\varepsilon_{0}<\frac{1}{4}\operatorname{Vol}(M^{n},g)$, by the Constancy Theorem it must be $\eta_{xy}=0$ and therefore $\tau(x)-\tau(y)=\tau_{xy}$ thus it is supported in $U_{C}$ as desired. Therefore $q(n-1,p,M,\delta)=1$, $K(N,n-1,p)=N$, $n(n-1,p)=0$, $\delta_{0}=\operatorname{inj}(M,g)$ and $c(n-1,p)=1$ for every $p\in\mathbb{N}$ and every $N\in\mathbb{N}.$
Inductive step 1. Assume that the result holds for families of $(k+1)$-cycles and let $F:X_{0}\to\mathcal{Z}_{k}(M)$ be as in the statement of the proposition. For each $x\in X_{0}$, we define $\tau(x)=\tau_{0}(x)$ to be an arbitrary $(k+1)$-chain such that $\partial\tau(x)=F(x)$ and $\operatorname{\mathbf{M}}(\tau(x))\leq\varepsilon$. Our goal is to define a sequence of maps $\tau_{j}:X_{j}(q_{j})_{0}\to\mathcal{I}_{k+1}(M)$ ($q_{j}=q(k,j,p,M,\delta)$ odd for each $j$) which agree with $\tau_{0}$ in $X_{0}$, are $\delta$-localized, small in mass and verify $\partial\tau_{j}(x)=R^{q_{j}}F(x)$. In order to do this, we consider the dyadic subdivision $X(2)$ of $X$ and for each $C\in\operatorname{Faces}_{p}(X)$ and $v\in V(C)$ we denote
@@
being $\{\overline{C}_{v}:C\in\operatorname{Faces}_{p}(X),v\in V(C)\}$ the collection of top dimensional cells of $X(2)$. Observe that by definition, if $x\in\overline{C}_{v}\cap X(q_{j})_{0}$ then $R^{q_{j}}F(x)=F(v)$ and hence $R^{q_{j}}F$ is constant on each $\overline{C}_{v}$. Therefore, in order to define $\tau_{j}$ inductively it will be easier to proceed one $\overline{C}_{v}$ at a time. This is because if $\tau_{j-1}$ was already defined in the boundary of a $j$-face $E$ of some $\overline{C}_{v}$ and $v_{E}$ is some vertex of $E$, then $\tilde{\tau}(x)=\tau_{j-1}(x)-\tau_{j-1}(v_{E})$ is a discrete family of $(k+1)$-cycles with domain $\partial E\cap X(q_{j-1})_{0}$ (observe that this is not true if $E$ is a cell of $C$ as the values of $\partial\tau_{j-1}(x)=R^{q_{j-1}}F(x)$ may vary). Therefore by the inductive hypothesis in the codimension of the cycles, there will exist $\sigma_{E}:\partial E\cap X(q_{j})_{0}\to\mathcal{I}_{k+2}(M)$ such that (up to refinement) $\partial\sigma_{E}(x)=\tilde{\tau}_{j}(x)$ and thus applying Lemma 3.23 to $\sigma_{E}$ we can extend $\sigma_{E}$ and therefore $\tau_{j-1}$ to all of $E$ as $\tau_{j}(x)=\tau_{j-1}(v_{E})+\partial\sigma_{E}(x)$. But we have to be careful with these extensions in order to obtain $\delta$-localized maps, because it may happen that two vertices $x,y$ are in the same cell of $X(q_{j})_{0}$ but do not belong to the same $\overline{C}_{v}$ (hence we can not define the $\tau_{j}$ on each $\overline{C}_{v}$ independently, some relations must hold for adjacent $\overline{C}_{v}$’s). We need to make sure that in that case the difference $\tau_{j}(x)-\tau_{j}(y)$ is still supported in a $\delta$-admissible family. For that purpose, we will consider a certain equivalence relation on the faces of the $\overline{C}_{v}$’s and extend $\tau_{j-1}$ one class at a time so that we maintain a certain relation between $\tau_{j}(x)$ and $\tau_{j}(y)$ if $x$ and $y$ are adjacent (namely, that the difference $\tau_{j}(x)-\tau_{j}(y)$ is supported in a $\delta$-admissible family).
@@
$\overline{C}_{v}$ =$=\{x\in C:d_{1}(x,v)\leq\frac{1}{2}\}$
$C_{v}$ =$=\{x\in C:d_{1}(x,v)\leq\frac{1}{3}\}$
@@
For every $q$, $\iota_{q}:\tilde{X}(q)\to X(2q+1)$ is a simplicial map. Moreover, its restriction to the 00-skeleton $\iota_{q}:\tilde{X}(q)_{0}\to X(2q+1)_{0}$ is bijective and is the identity when $q=1$. Therefore, there is a bijective correspondence between discrete maps $\tilde{F}:\tilde{X}(q)_{0}\to\mathcal{I}_{k}(M)$ and discrete maps $F:X(2q+1)_{0}\to\mathcal{I}_{k}(M)$ given by by $F\circ\iota_{q}=\tilde{F}$. We will often refer to $F$ and $\tilde{F}$ using the same symbol considering this identification.
@@
Let $\sim$ be the equivalence relation on the cells of $\tilde{X}$ given by $E\sim E^{\prime}$ if and only if $\iota(E)=\iota(E^{\prime})$. If $E\sim E^{\prime}$, we define an affine map $T_{EE^{\prime}}:E\to E^{\prime}$ given by $T_{EE^{\prime}}=\iota|_{E^{\prime}}^{-1}\circ\iota|_{E}$. Given $F,F^{\prime}\in\operatorname{Faces}_{j}(\tilde{X}(q_{j}))$ we say that $F\sim F^{\prime}$ if there exist $E,E^{\prime}\in\operatorname{Faces}_{j}(X)$ such that $F\subseteq E$, $F^{\prime}\subseteq E^{\prime}$, $E\sim E^{\prime}$ and $T_{EE^{\prime}}(F)=F^{\prime}$. In that case, we denote $T_{FF^{\prime}}=T_{EE^{\prime}}$
@@
Taking the previous observation in mind, we are going to inductively define $\delta_{j}$-localized maps $\tau_{j}:\tilde{X}_{j}(q_{j})_{0}\to\mathcal{I}_{k+1}(M)$ satisfying the following: if $F\sim F^{\prime}$, $F,F^{\prime}\in\operatorname{Faces}_{j}(\tilde{X}(q_{j}))$ and $x\in F$, the chain $\Delta^{j}_{FF^{\prime}}(x):=\tau_{j}(x)-\tau_{j}(T_{FF^{\prime}}(x))$ is supported in a certain $\delta_{j}$-admissible family $\{B_{i}^{F}\}_{i\in I_{F}}$. This will allow us to induce a map $\tau$ on $X(2q_{p}+1)_{0}$ from $\tau_{p}$ which will be $\delta$-localized and will have the other desired properties.

### v2

@@ 3.2. Main Theorems
In this subsection, we state the approximation results by $\delta$-localized families in the discrete and continuous setting. These are crucial in the proof of the Parametric Isoperimetric Inequality discussed later in this article and also to obtain the Parametric Coarea Inequality in [StaCoarea]. For simplicity, they are stated for families of absolute cycles and chains, but the corresponding relative versions in $(M,\partial M)$ also hold. They apply to any coefficient group $\mathbb{G}$ ($\mathbb{G}=\mathbb{Z}$ or $\mathbb{G}=\mathbb{Z}_{p}$). The manifold $M$ is compact, of dimension at least $2$.
@@
Given $0\leq k\leq n-1$ and $p,N\in\mathbb{N}$ there exist natural numbers $K(N,k,p)$, $n(k,p)$ and positive constants $c(k,p,M)$, $\varepsilon_{0}(k,p,M)$ and $\delta_{0}(k,p,M)$ so that the following is true. Let $\varepsilon\leq\varepsilon_{0}$, $\delta\leq\delta_{0}$ and let $F:X_{0}\to\mathcal{Z}_{k}(M)$ be a discrete $(N,\delta)$-localized family of cycles with $\mathcal{F}(F(x))\leq\varepsilon$ for every $x\in X_{0}$ and $\dim(X)=p$. Then there exists a $(K(N,k,p),K(N,k,p)\delta)$-localized family $\tau:X(q)_{0}\to\mathcal{I}_{k+1}(M)$ such that $\partial\tau(x)=R^{q}F(x)$ for $q=q(k,p,M,\delta)$ and $\operatorname{\mathbf{M}}(\tau(x))\leq c(k,p,M)\frac{\varepsilon}{\delta^{n(k,p)}}$.
Roughly speaking, the previous theorem tells us that if we have a discrete family of $k$-cycles $F:X_{0}\to\mathcal{Z}_{k}(M)$ which has flat norm at most $\varepsilon$ at every point and is $\delta$-localized for $\varepsilon$ and $\delta$ small enough, then we can construct a discrete filling with similar properties. To be precise, the filling is a $\delta^{\prime}$-localized family of $(k+1)$-chains $\tau$ such that up to refinement $\partial\tau(x)=F(x)$, $\delta^{\prime}$ is controlled in terms of $\delta$ and $\operatorname{\mathbf{M}}(\tau(x))$ is controlled in terms of $\varepsilon$ and $\delta$. Theorem 3.9 is proved in subsection 3.3 and is used to obtain the following result.
@@
There exist constants $\varepsilon_{0}(k,p,M)>0$, $\delta_{0}(k,p,M)>0$, $C(k,p,M)>0$, $n(k,p)\in\mathbb{N}$ and $N(k,p)\in\mathbb{N}$ such that the following is true. Let $F:X_{0}\to\mathcal{Z}_{k}(M)$ be an $\varepsilon$-fine family, $0<\varepsilon\leq\varepsilon_{0}$, $0\leq k\leq n-1$, $\dim(X)=p$. Let $0<\delta<\delta_{0}$. Then there exists a discrete family $F^{\prime}:X(q)_{0}\to\mathcal{Z}_{k}(M)$ for $q=q(k,p,M,\delta)$ such that
@@
There exist constants $\varepsilon_{0}(k,p,M)>0$, $\delta_{0}(k,p,M)>0$, $C(k,p,M)$, $n(k,p)\in\mathbb{N}$ and $N(k,p)\in\mathbb{N}$ such that the following is true. Let $G:X_{0}\to\mathcal{I}_{k+1}(M)$ be an $\varepsilon$-fine family, $0<\varepsilon\leq\varepsilon_{0}$, $0\leq k\leq n-1$, $\dim(X)=p$ and denote $F=\partial G$. Let $0<\delta<\delta_{0}$. Then there exists a discrete family $G^{\prime}:X(q)_{0}\to\mathcal{I}_{k+1}(M)$ for $q=q(k,p,M,\delta)$ such that if we denote $F^{\prime}=\partial G^{\prime}$ then
@@
Let $G:X\to\mathcal{I}_{k+1}(M)$ be a continuous map without concentration of mass. Denote $F=\partial G$ and let $\varepsilon,\delta>0$. Then there exists $q,q^{\prime}\in\mathbb{N}$ and a continuous map $G^{\prime}:X\to\mathcal{I}_{k+1}(M)$ without concentration of mass such that if we denote $F^{\prime}=\partial G^{\prime}$ then
@@
3.3. Filling discrete families of small cycles
@@
Given $0\leq k\leq n-1$ and $p,N\in\mathbb{N}$ there exist natural numbers $K(N,k,p)$, $n(k,p)$ and positive constants $c(k,p)$, $\varepsilon_{0}(k,p,M)$ and $\delta_{0}(k,p,M)$ so that the following is true. Let $\varepsilon\leq\varepsilon_{0}$, $\delta\leq\delta_{0}$ and let $F:X_{0}\to\mathcal{Z}_{k}(M)$ be a discrete $(N,\delta)$-localized family of cycles with $\mathcal{F}(F(x))\leq\varepsilon$ for every $x\in X_{0}$ and $\dim(X)=p$. Then there exists a $(K(N,k,p),K(N,k,p)\delta)$-localized family $\tau:X(q)_{0}\to\mathcal{I}_{k+1}(M)$ such that $\partial\tau(x)=R^{q}F(x)$ for $q=q(k,p,M,\delta)$ and $\operatorname{\mathbf{M}}(\tau(x))\leq c(k,p)(\frac{L}{\delta})^{n(k,p)}\varepsilon$.
In order to prove the proposition, we need to introduce some definitions and do some constructions first. Let $r<\operatorname{inj}(M,g)$ and consider a collection of $L$ points $x_{1},...,x_{L}\in M$ such that $\{B(x_{l},r):1\leq l\leq L\}$ covers $M$ and $L\leq c(n)\frac{\operatorname{Vol}(M,g)}{r^{n}}$ for some universal constant $c(n)$ depending only on $n$. For this section, we will consider $r=\delta$ and the corresponding $L$ is the one that appears in the statement of Proposition 3.14. In the next sections, we will choose other values of $r$, being that the reason why we do the following constructions for arbitrary $r<\operatorname{inj}(M,g)$.
@@
$\operatorname{\mathbf{M}}(\tau\llcorner\partial B(x_{l},r_{l}(\tau)))\leq\frac{\operatorname{\mathbf{M}}(\tau)}{r}.$
@@
We introduce the following definitions, which will be used in the rest of the section.
@@
$d^{q}(x,y)$ equals the minimal number of vertices in $X(q)_{0}$ that we must visit if we want to travel from $x$ to $y$ by jumping between vertices in $X(q)_{0}$ differing in only one coordinate. When $q$ is clear (because we are considering $x,y\in X(q)_{0}$ for some specific value of $q$) we may just denote $d^{q}$ by $d$.
Later we will be interested in constructing discrete $\delta$-localized families inductively skeleton by skeleton on a cubical complex $X$. When doing that, we will have to extend a family $\tau(x)$ defined on the boundary of a certain cell $C$ of $X$ to the interior of $C$, preserving $\delta$-localization. The next lemma shows that this is possible if the extension is done by a chopping procedure as in Definition 3.16. The subsequent lemma shows that if in addition $\operatorname{\mathbf{M}}(\tau(x))$ and $\operatorname{\mathbf{M}}(\partial\tau(x))$ are small, then the extension and its boundary also have small mass.
@@
$d_{l}(\tau)-d_{l}(\tau^{\prime})=\tau-\tau^{\prime}-[\tau\llcorner\bigcup_{m=1}^{l}B(x_{m},r_{m})-\tau^{\prime}\llcorner\bigcup_{m=1}^{l}B(x_{m},r^{\prime}_{m})]\\ =\tau-\tau^{\prime}-[(\tau-\tau^{\prime})\llcorner\bigcup_{m=1}^{l}B(x_{m},r_{m}^{\prime})+\tau\llcorner\bigcup_{m=1}^{l}B(x_{m},r_{m})\setminus B(x_{m},r^{\prime}_{m})\\ -\tau\llcorner\bigcup_{m=1}^{l}B(x_{m},r^{\prime}_{m})\setminus B(x_{m},r_{m})].$
The previous implies that
@@
First we prove $(N,\delta)$-localization. Let $y_{1},y_{2}$ be vertices of the same cell $E$ in $C(3q)$. If $y_{2},y_{3}\in\operatorname{Center}(C)$ the result follows, so we can assume $l(y_{1}),l(y_{2})\leq q$. Let $E^{\prime}$ be a (j1)(j-1)-face of $\partial C(q)$ such that $\Xi(E)\subseteq E^{\prime}$. Then without loss of generality, there exists another vertex $y_{3}$ of $E$ such that $\Xi(y_{2})=\Xi(y_{3})$, $l(y_{2})=l(y_{3})+1$, $l(y_{1})=l(y_{3})$ and $\Xi(y_{1}),\Xi(y_{3})$ belong to $E^{\prime}$. By Lemma 3.19 applied to the chains $\tau=\tau(\Xi(y_{1}))$, $\tau^{\prime}=\tau(\Xi(y_{3}))$, the number $l=l(y_{1})=l(y_{3})$ and the $(N,\delta)$-admissible family $\{U_{i}\}_{i\in I_{E^{\prime}}}$ with $U_{i}=B(y_{i},s_{i})$ it holds
@@
For each cell $C$ of $X$, let $\{U_{i}^{C}\}_{i\in I_{C}}$ be the $\delta$-admissible family associated to $C$ and let $U_{C}=\bigcup_{i\in I_{C}}U_{i}^{C}$. We proceed by induction on nkn-k, being k=n1k=n-1 the base case.
Base case 1. Let k=n1k=n-1. For each $x\in X_{0}$, as $\mathcal{F}(F(x))\leq\varepsilon$ there exists an $n$-chain $\tau(x)$ such that $\partial\tau(x)=F(x)$ and $\operatorname{\mathbf{M}}(\tau(x))\leq\varepsilon$. Consider the corresponding family $\tau:X_{0}\to\mathcal{I}_{n}(M)$. We claim that if $C$ is a cell of $X$ and $x,y\in X_{0}\cap C$ then $\tau(x)-\tau(y)$ is supported in $U_{C}$. Indeed, as $\operatorname{spt}(F(x)-F(y))\subseteq U_{C}$, there exists $\tau_{xy}\in\mathcal{I}_{n}(U_{C})$ such that $\partial\tau_{xy}=F(x)-F(y)$ and $\operatorname{\mathbf{M}}(\tau_{xy})\leq 2\varepsilon$. Therefore $\eta_{xy}=\tau(x)-\tau(y)-\tau_{xy}$ is an $n$-cycle on $M$ of mass at most $4\varepsilon$. Thus if we set $\varepsilon_{0}<\frac{1}{4}\operatorname{Vol}(M^{n},g)$, by the Constancy Theorem it must be $\eta_{xy}=0$ and therefore $\tau(x)-\tau(y)=\tau_{xy}$ thus it is supported in $U_{C}$ as desired. Therefore $q(n-1,p,M,\delta)=1$, K(N,n1,p)=NK(N,n-1,p)=N, n(n1,p)=0n(n-1,p)=0, $\delta_{0}=\operatorname{inj}(M,g)$ and c(n1,p)=1c(n-1,p)=1 for every $p\in\mathbb{N}$ and every $N\in\mathbb{N}.$
Inductive step 1. Assume that the result holds for families of $(k+1)$-cycles and let $F:X_{0}\to\mathcal{Z}_{k}(M)$ be as in the statement of the proposition. For each $x\in X_{0}$, we define $\tau(x)=\tau_{0}(x)$ to be an arbitrary $(k+1)$-chain such that $\partial\tau(x)=F(x)$ and $\operatorname{\mathbf{M}}(\tau(x))\leq\varepsilon$. Following the strategy of Guth and Liokumovich [GL22], we define a sequence of maps $\tau_{j}:X_{j}(q_{j})_{0}\to\mathcal{I}_{k+1}(M)$ ($q_{j}=q(k,j,p,M,\delta)$ odd for each $j$) which agree with $\tau_{0}$ at $X_{0}$, are $\delta$-localized, small in mass and verify $\partial\tau_{j}(x)=R^{q_{j}}F(x)$. In order to do this, we consider the dyadic subdivision $X(2)$ of $X$ and for each $C\in\operatorname{Faces}_{p}(X)$ and $v\in V(C)$ we denote
@@
being $\{\overline{C}_{v}:C\in\operatorname{Faces}_{p}(X),v\in V(C)\}$ the collection of top dimensional cells of $X(2)$. Observe that by definition, if $x\in\overline{C}_{v}\cap X(q_{j})_{0}$ then $R^{q_{j}}F(x)=F(v)$ and hence $R^{q_{j}}F$ is constant on each $\overline{C}_{v}$. Therefore, in order to define $\tau_{j}$ inductively it will be easier to proceed one $\overline{C}_{v}$ at a time. This is because if $\tau_{j-1}$ was already defined in the boundary of a $j$-face $E$ of some $\overline{C}_{v}$ and $v_{E}$ is some vertex of $E$, then $\tilde{\tau}(x)=\tau_{j-1}(x)-\tau_{j-1}(v_{E})$ is a discrete family of $(k+1)$-cycles with domain $\partial E\cap X(q_{j-1})_{0}$ (observe that this is not true if $E$ is a cell of $C$ as the values of $\partial\tau_{j-1}(x)=R^{q_{j-1}}F(x)$ may vary). Therefore by the inductive hypothesis on the codimension of the cycles, there will exist $\sigma_{E}:\partial E\cap X(q_{j})_{0}\to\mathcal{I}_{k+2}(M)$ such that (up to refinement) $\partial\sigma_{E}(x)=\tilde{\tau}_{j}(x)$ and thus applying Lemma 3.20 to $\sigma_{E}$ we can extend $\sigma_{E}$ and therefore $\tau_{j-1}$ to all of $E$ as $\tau_{j}(x)=\tau_{j-1}(v_{E})+\partial\sigma_{E}(x)$. But we have to be careful with these extensions in order to obtain $\delta$-localized maps, because it may happen that two vertices $x,y$ are in the same cell of $X(q_{j})_{0}$ but do not belong to the same $\overline{C}_{v}$ (hence we can not define the $\tau_{j}$ on each $\overline{C}_{v}$ independently, some relations must hold for adjacent $\overline{C}_{v}$’s). We need to make sure that in that case the difference $\tau_{j}(x)-\tau_{j}(y)$ is still supported in a $\delta$-admissible family. For that purpose, we will consider a certain equivalence relation on the faces of the $\overline{C}_{v}$’s and extend $\tau_{j-1}$ one class at a time so that we maintain a certain relation between $\tau_{j}(x)$ and $\tau_{j}(y)$ if $x$ and $y$ are adjacent (namely, that the difference $\tau_{j}(x)-\tau_{j}(y)$ is supported in a $\delta$-admissible family).
@@
$\overline{C}_{v}$ $=\{x\in C:d_{1}(x,v)\leq\frac{1}{2}\}$
$C_{v}$ $=\{x\in C:d_{1}(x,v)\leq\frac{1}{3}\}$
@@
For every $q$, $\iota_{q}:\tilde{X}(q)\to X(2q+1)$ is a simplicial map. Moreover, its restriction to the 0-skeleton $\iota_{q}:\tilde{X}(q)_{0}\to X(2q+1)_{0}$ is bijective and is the identity when $q=1$. Therefore, there is a bijective correspondence between discrete maps $\tilde{F}:\tilde{X}(q)_{0}\to\mathcal{I}_{k}(M)$ and discrete maps $F:X(2q+1)_{0}\to\mathcal{I}_{k}(M)$ given by by $F\circ\iota_{q}=\tilde{F}$. We will often refer to $F$ and $\tilde{F}$ using the same symbol considering this identification.
@@
Let $\sim$ be the equivalence relation on the cells of $\tilde{X}$ given by $E\sim E^{\prime}$ if and only if $\iota(E)=\iota(E^{\prime})$. If $E\sim E^{\prime}$, we define an affine map $T_{EE^{\prime}}:E\to E^{\prime}$ given by $T_{EE^{\prime}}=\iota|_{E^{\prime}}^{-1}\circ\iota|_{E}$. Given $F,F^{\prime}\in\operatorname{Faces}_{j}(\tilde{X}(q_{j}))$ we say that $F\sim F^{\prime}$ if there exist $E,E^{\prime}\in\operatorname{Faces}_{j}(X)$ such that $F\subseteq E$, $F^{\prime}\subseteq E^{\prime}$, $E\sim E^{\prime}$ and $T_{EE^{\prime}}(F)=F^{\prime}$. In that case, we denote $T_{FF^{\prime}}=T_{EE^{\prime}}$.
@@
[…]